# Few-shot Species Range Estimation

### v1

@@ A.1.2 Training
For all training we use the Adam optimizer (Kingma & Ba, 2015) with a learning rate of 0.0005, and an exponential learning rate scheduler with a learning rate decay of 0.98 per epoch, and we use a batch size of 2048. Our training data comes from Cole et al. (2023), comprising of 35.5 million species observations with locations, covering 47,375 species observed prior to 2022 on the iNaturalist platform. However, we remove all species that are found in our evaluation datasets, leaving us with 44,181 species in our training set.
@@
Training comprises of two steps. First, the location encoder, $f$, is trained. This follows the training procedure of Cole et al. (2023) using the $\mathcal{L}_{\text{AN-full}}$ loss function with the positive weighting, $\lambda$, set to 2,048, training for 20 epochs with a dropout of 0.5. To reduce training time without significantly impacting performance we only train on a maximum of 1,000 examples per-species, as done in Cole et al. (2023). Thus our training dataset for this step contains 13.8 million location observations. Secondly, we train all components of FS-SINR, except the pretrained large language model and the pretrained vision transformer, using our $\mathcal{L}_{\text{AN-full-b}}$ loss with $\lambda$ set to 2,048. We train the location encoder, $f$, again as this improves performance compared to freezing it, seen in Figure A5. For this part of training we use a dropout of 0.2. We further reduce the training data used to a maximum of 100 examples per-species, leaving 4.0 million training examples, which again increases training speed without a significant impact on performance, as seen in Figure A6. For this step we additionally train with images and text descriptions of the training species. Each instance in the training set is used once per epoch as a training example to compute the loss. The training example is not passed through the transformer encoder, $e$, and so does not contribute to making the species embedding vector produced by this part of the model. Instead, additional context information is provided to produce the species embedding. By default this context information is comprised of 20 context locations, a section of text describing the target species and an image of the species. With 0.2 probability the context locations are dropped from the context information, and with 0.5 probability each the text or image is dropped. If all modalities are dropped we skip that training example. These context locations are taken from the training data for the target species. As such, a single instance from the training set can be used multiple times per epoch, once as a training example, and potentially many times as a context location. The impact of different distributions of locations and text provided during training is shown in Figure A3.
@@
For the text inputs required during this stage of training, we use the text dataset from Hamilton et al. (2024) comprising of multiple sections of Wikipedia (Wikipedia, 2025) articles for each species in the train set where these are available. This dataset contains 127,484 sections from 37,889 species’ articles. The evaluation text either describes the habitat or range a species, where habitat text tends to describe the local environment and range text is typically more informative as it often lists specific countries or geographic regions where the species can be found. Note, that not all 44,181 train species have text data available. The images used are taken from iNaturalist (2025), and this dataset comprises 204,064 images of our train species. When an image or piece text is not available for a species during training, and we are trying to provide these modalities and context locations to the model, we simply ignore the additional modality and only provide the context locations. When we are attempting to provide just an image or text as context, we instead skip that training instance.
@@
In practice, during training, we pass all text sections through the frozen large language model once and then store the embeddings produced to use in the current training run and all future runs, and similarly extract and store all image embeddings after passing the images through the frozen vision transformer. This prevents us having to repeatedly query these frozen, but resource intensive, models during training. Training takes approximately ten hours on a single NVIDIA A6000 GPU, requiring approximately six gigabytes of RAM.
A.2 Baselines
A.2.1 LE-SINR
@@
We also follow the original evaluation procedure for LE-SINR. For few-shot evaluation without text, logistic regression with L2 regularization is performed with location features as input using the few positive examples provided alongside a set of pseudo-negatives drawn half from a uniform random distribution and half from the training data distribution. The regularization weight is set to 20. For text-based zero-shot evaluation we directly make use of the output of the text encoder with the dot product between this and location features giving us a probability of species presence. For few-shot evaluation, when text is provided, we again perform logistic regression, but the output of the text encoder is used as the “target” that the weights are drawn towards in a modified L2 regularization term – see Hamilton et al. (2024) for more details. The regularization weight is again set to 20. In total this model comprises 25,715,202 learnable parameters.
A.2.2 SINR
We also compare to SINR (Cole et al., 2023). The original SINR implementation requires all evaluation species to be part of the training set. We match the adaptations from Hamilton et al. (2024) to allow evaluation on unseen species. After training we remove the learned species heads and keep only the location encoder. During evaluation we perform logistic regression with L2 regularization using location features as input. The regularization weight is again set to 20, and the same method of selecting pseudo-negatives as above is used. In total this model comprises 11,941,120 learnable parameters.
A.2.3 Prototype SINR
Here we describe few-shot baselines based on Prototypical Networks (Snell et al., 2017). Our approach is very similar to Snell et al. (2017) although we use the SINR location encoder of our models as the ‘embedding function’, allowing us to generate few-shot results for a novel species without any retraining. This SINR location encoder is trained only on species found in the training set and not those used for evaluation. Using this method, SINR and LE-SINR models can be used to estimate the range of a novel species without requiring training to learn a new species vector.
@@ A.2.3 Prototype SINR
where $k\in\{\text{present},\text{absent}\}$ indicates the class of the prototype, and $S$ is the ‘support set’, i.e., the set of locations $\bm{x}$ that we use to create our prototypes. In our case, $S_{present}$ is the small set of available context locations for our target species, i.e., $\mathcal{C}^{t}$, while $S_{absent}$ is the set of pseudo-negative locations that we have selected according to Hamilton et al. (2024). $f_{\bm{\theta}}()$ indicates the location encoder of our model.
To generate a probability of presence or absence at any location $\bm{x}$, we encode $\bm{x}$ using our location encoder and calculate the cosine distance in ‘location encoder space’ between $\bm{x}$ and each prototype. We then use these values as the ‘logits’ in a softmax function to generate our probabilities. The parameters of the location encoder are not changed. Putting this together, we can calculate the probability of presence as:
@@
While the original implementation in Snell et al. (2017) uses the squared Euclidean distance instead of cosine distance we find that this performs significantly worse and actually results in decreased MAP as the number of context locations increases. We suggest the SINR location encoder is more suited to using cosine distance as during training, presence predictions are generated by taking the dot product of the location and species embeddings. However when the location encoder is trained from scratch for prototype SINR as in Snell et al. (2017) we find that using the squared Euclidean distance performs better than using cosine distance, though performance is still lower than cosine distance with a SINR location encoder.
@@
In Figure 3 we see that the performance of the ‘prototype’ approach is worse than FS-SINR and the SINR and LE-SINR baselines. In Table A1 we provide zero-shot results where the species embedding produced from text as with LE-SINR zero-shot predictions is used as a positive prototype. In both cases, prototype SINR underperforms compared to our approach. In Figure A1 we present qualitative results visualizing the few-shot estimated range for the Kalahari Scrub-Robin produced by FS-SINR and by the Prototype SINR baseline.
@@
Figure A1: Qualitative comparison of the Prototype SINR baseline. Here we compare the predictions of our FS-SINR approach (without any text) and our ‘Prototype SINR’ baseline on the Kalahari Scrub-Robin species that is found in Southern Africa. The Prototype SINR approach obtains an MAP of 0.54 and 0.79, for one and two context locations respectively, while FS-SINR obtains 0.79 and 0.85. As MAP is tied to the ranking of predicted probabilities rather than their absolute values, it can remain high even if the model is somewhat overconfident across the board. As long as the highest probabilities consistently align with areas where the species is truly present, the model will achieve a strong MAP, which we can see with the predictions from Prototype SINR.
A.2.4 Active SINR
@@
We also compare to the model introduced for active learning in Lange et al. (2023), which we call Active SINR, though in our setting there is no active learning component. This approach begins with a SINR model trained on our reduced 44,181 species which do not include the evaluation species. The weights $\bm{W}$ of the multi-label classifier of this model can be viewed as a set of species embeddings where each column vector $\bm{w_{j}}$ of $\bm{W}$ represents an individual species $j$. We can combine these species embeddings with a location embedding $f_{\bm{\theta}}(\bm{x})$ via an inner product to compute the probability that the species $j$ is present at $\bm{x}$. At inference time, we compute the presence probabilities for all species in the training set, for all locations $c$ in the set of available context locations $\mathcal{C}^{t}$ for our target species $t$. We then produce a new species embedding $\bm{w}_{t}$ by taking a weighted average of the existing $\bm{w}_{j}$’s where the weight for each is the product of the probabilities of presence for that species:
@@
We can then use this new species embedding for our target species to produce a probability of presence for any location $x$ as in SINR (Cole et al., 2023). We present few-shot results using this method in Figure 3. We see that the performance of the Active SINR approach is competitive with FS-SINR when provided with no text, though worse than FS-SINR when provided with this additional context. However increasing the number of provided context locations beyond a small number actually hurts performance, as it is unable to accurately represent the range of a previously unseen species via the weighted combination of those from the training set.
A.3 Evaluation
@@ Appendix B Ablations
Here we present additional results to investigate the impact of a variety of design choices and training procedures for FS-SINR. We present plots on a “Symlog” scale, where a linear scale is used between 0 and 1 in order to allow us to show zero-shot results alongside few-shot results. We display the mean of three runs with standard deviations shown as error bars and also show just the mean values alongside for easier interpretation.
B.1 Ablating Training Context Locations
@@
In Figure A2 we show “Range Text” evaluation performance on the IUCN dataset for FS-SINR models trained using different amounts of context information at training time. We see that generally increasing the context used during training improves performance, and that having a fixed number of context locations is also beneficial.
@@
Figure A2: Impact of number of training context locations. Here we evaluate FS-SINR models trained using different numbers of context locations. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with “Range Text” on the IUCN dataset. “Fixed” indicates the same number of context locations were provided for every training example. “Variable” indicates that a uniform random distribution of context locations up to the specified number were provided with each training example. We see that “Variable” generally under-performs compared to “Fixed” and that increasing the train context length tends to increase evaluation performance.
B.2 Ablating Context Information
@@
In Figure A3 we display “Range Text” evaluation performance on the IUCN dataset for FS-SINR models trained using different combinations of text and location context information during training. We observe that good text-only zero-shot performance requires sometimes providing just text as context information during training. This forces the model to learn to produce ranges from only text information. Models that are sometimes provided with both text and locations for the same training examples perform best as the number of provided context locations increases. We also see that models trained without text can perform on par with those that see text during training when enough context locations are provided (5 - 10). As we might expect, models that are provided with token types they have not seen during training perform poorly.
@@
Figure A3: Impact of train context information. Here we evaluate FS-SINR models trained using different context information on the IUCN dataset. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with “Range Text” unless “No Eval Text” is specified, in which case just locations are provided during eval. 70% of training examples for “Default FS-SINR” provide both location and text context, 20% provide just locations 10% and provide just text. “Always Obs. Only” has only seen locations during training. “Always Text Only” has only seen Text during training. “Always Text and Obs” is always provided with both locations and text during training. “Always Text or Obs.” is provided with just locations for 90% of training examples, and just text for the remaining 10%.
B.3 Ablating Input Features
@@
In Table A1 we provide additional zero-shot results expanding on those in Table 1 from the main paper. Specifically, we add comparisons to using a different location encoder (i.e., SATCLIP (Klemmer et al., 2023) instead of SINR), comparisons to using a DINOv2 pretrained image encoder (DINOv2-large), comparisons to using the environmental covariates as in SINR (Cole et al., 2023) that contain information about a locations’ climate in addition to the spatial coordinates.
@@
Table A1: Additional zero-shot results. We report zero-shot performance where no location information is provided to each model, comparing SINR (Cole et al., 2023), LE-SINR (Hamilton et al., 2024), and variants. We denote additional metadata as: EN for additional environmental covariates (Fick & Hijmans, 2017) used in Cole et al. (2023), HT for “Habitat Text”, RT for “Range Text”, I for “Image”, DINOV2 for Image using a DINOV2 based image encoder (Oquab et al., 2024), TST for “Test Species in Train”, TRT for using full taxonomic rank text, SATCLIP for where the SINR encoders are replaced with the image derived location encoders from Klemmer et al. (2023), and P-LE-SINR for “Prototype LE-SINR”. Results are reported as MAP, where higher is better.
@@
B.4 Ablating Location Encoder

### v2

@@ C.1.2 Training
For all training we use the Adam optimizer (Kingma & Ba, 2015) with a learning rate of 0.0005, and an exponential learning rate scheduler with a learning rate decay of 0.98 per epoch, and we use a batch size of 2048. Our training data comes from Cole et al. (2023), comprising 35.5 million species observations with locations, covering 47,375 species observed prior to 2022 on the iNaturalist platform. However, we remove all species found in our evaluation datasets, leaving us with 44,181 species in our training set.
@@
Training comprises of two steps. First, the location encoder, $f$, is trained. This follows the training procedure of Cole et al. (2023) using the $\mathcal{L}_{\text{AN-full}}$ loss function with positive weighting, $\lambda$, set to 2,048, training for 20 epochs with a dropout of 0.5. To reduce training time without significantly impacting performance we only train on a maximum of 1,000 examples per-species, as done in Cole et al. (2023). Thus, our training dataset for this step contains 13.8 million location observations. Secondly, we train all components of FS-SINR, except the pre-trained large language model and the pretrained vision transformer, using our $\mathcal{L}_{\text{AN-full-b}}$ loss with $\lambda$ set to 2,048. We train the location encoder, $f$, again as this improves performance compared to freezing it,as seen in Figure A15. For this part of training, we use a dropout of 0.2. We further reduce the training data used to a maximum of 100 examples per-species, leaving 4.0 million training examples, which again increases training speed without a significant impact on performance, as seen in Figure A16. For this step we additionally train with images and text descriptions of the training species. Each instance in the training set is used once per epoch as a training example to compute the loss. The training example is not passed through the transformer encoder, $m$, and so does not contribute to the species embedding vector produced by this part of the model. Instead, additional context information is provided to produce the species embedding. By default, this context information consists of 20 context locations, a section of text describing the target species, and an image of the species. With 0.2 probability the context locations are dropped from the context information, and with 0.5 probability each the text or image is dropped. These context locations are taken from the training data for the target species. As such, a single instance from the training set can be used multiple times per epoch, once as a training example, and potentially many times as a context location. The impact of different distributions of locations and text provided during training is shown in Figure A13.
@@
For the text inputs required during this stage of training, we use the text dataset from Hamilton et al. (2024) consisting of multiple sections of Wikipedia (Wikipedia, 2025) articles for each species in the train set where these are available. This dataset contains 127,484 sections from 37,889 species’ articles. The evaluation text either describes the habitat or range a species, where habitat text tends to describe the local environment and range text is typically more informative as it often lists specific countries or geographic regions where the species can be found. Note that not all 44,181 train species have text data available. The images used are taken from iNaturalist (2025), and this dataset comprises 204,064 images of our train species. When an image or piece of text is not available for a species during training, and we are trying to provide these modalities and context locations to the model, we simply ignore the additional modality and only provide the context locations. When we are attempting to provide just an image or text as context, we instead skip that training instance.
@@
In practice, during training, we pass all text sections through the frozen large language model once and then store the embeddings produced to use in the current training run and all future runs, and similarly extract and store all image embeddings after passing the images through the frozen vision transformer. This prevents us having to repeatedly query these frozen, but resource-intensive, models during training. Training takes approximately ten hours on a single NVIDIA A6000 GPU, requiring approximately six gigabytes of RAM.
C.2 Baselines
C.2.1 LE-SINR
@@
We also follow the original evaluation procedure for LE-SINR. For few-shot evaluation without text, logistic regression with L2 regularization is performed with location features as input using the few positive examples provided alongside a set of pseudo-negatives drawn half from a uniform random distribution and half from the training data distribution. The regularization weight is set to 20. For text-based zero-shot evaluation, we directly make use of the output of the text encoder with the dot product between this and location features giving us a probability of species presence. For few-shot evaluation, when text is provided, we again perform logistic regression, but the output of the text encoder is used as the ‘target’ that the weights are drawn towards in a modified L2 regularization term,see Hamilton et al. (2024) for more details. The regularization weight is again set to 20. In total, this model comprises 25,715,202 learnable parameters.
C.2.2 SINR
We also compare to SINR (Cole et al., 2023). The original SINR implementation requires all evaluation species to be part of the training set. We match the adaptations from Hamilton et al. (2024) to allow evaluation on unseen species. After training we remove the learned species heads and keep only the location encoder. During evaluation, we perform logistic regression with L2 regularization using location features as input. The regularization weight is again set to 20, and the same method of selecting pseudo-negatives as above is used. In total, this model comprises 11,941,120 learnable parameters.
C.2.3 Prototype SINR
Here we describe our few-shot baseline based on Prototypical Networks (Snell et al., 2017), which we refer to as Prototype SINR. Our approach is very similar to Snell et al. (2017) although we use the SINR location encoder of our models as the ‘embedding function’, allowing us to generate few-shot results for a novel species without any retraining. This SINR location encoder is trained only on species found in the training set and not those used for evaluation. Using this method, SINR and LE-SINR models can be used to estimate the range of a novel species without requiring training to learn a new species embedding vector.
@@ C.2.3 Prototype SINR
where $k\in\{\text{present},\text{absent}\}$ indicates the class of the prototype, and $S$ is the ‘support set’, i.e., the set of locations $\bm{x}$ that we use to create our prototypes. In our case, $S_{present}$ is the small set of available context locations for our target species, i.e., $\mathcal{C}^{t}$, while $S_{absent}$ is the set of pseudo-negative locations that we have selected according to Hamilton et al. (2024). $f_{\bm{\theta}}()$ denotes the location encoder of our model.
To generate a probability of presence or absence for any location $\bm{x}$, we encode $\bm{x}$ using our location encoder and calculate the cosine distance in ‘location encoder space’ between $\bm{x}$ and each prototype. We then use these values as the ‘logits’ in a softmax function to generate our probabilities. The parameters of the location encoder are not changed. Putting this together, we can calculate the probability of presence as:
@@
While the original implementation in Snell et al. (2017) uses the squared Euclidean distance instead of cosine distance we find that this performs significantly worse and actually results in decreased MAP as the number of context locations increases. We suggest the SINR location encoder is more suited to using cosine distance, as during training presence predictions are generated by taking the dot product of the location and species embeddings. However, when the location encoder is trained from scratch for Prototype SINR as in Snell et al. (2017) we find that using the squared Euclidean distance performs better than cosine distance, although performance is still lower than cosine distance with a SINR location encoder.
@@
In Figure 3 we see that the performance of Prototype SINR is worse than FS-SINR and the SINR and LE-SINR baselines. In Table A5 we provide zero-shot results where the positive prototype is a species embedding produced from text, in the same manner as LE-SINR zero-shot predictions. In both cases, Prototype SINR underperforms compared to our approach. In Figure A11 we present qualitative results visualizing the few-shot estimated range for the Kalahari Scrub-Robin produced by FS-SINR and by the Prototype SINR baseline.
@@
Figure A11: Qualitative comparison of the Prototype SINR baseline. Here we compare the predictions of our FS-SINR approach (without any text) and our Prototype SINR baseline on the Kalahari Scrub-Robin species that is found in Southern Africa. The Prototype SINR approach obtains an MAP of 0.54 and 0.79, for one and two context locations respectively, while FS-SINR obtains 0.79 and 0.85. As MAP is tied to the ranking of predicted probabilities rather than their absolute values, it can remain high even if the model is somewhat overconfident across the board. As long as the highest probabilities consistently align with areas where the species is truly present, the model will achieve a strong MAP, which we can see with the predictions from Prototype SINR.
C.2.4 Active SINR
@@
We also compare to the model introduced for active learning in Lange et al. (2023), which we call Active SINR, although in our setting there is no active learning component. This approach begins with a SINR model trained on our reduced 44,181 species which do not include the evaluation species. The weights $\bm{W}$ of the multi-label classifier of this model can be viewed as a set of species embeddings where each column vector $\bm{w_{j}}$ of $\bm{W}$ represents an individual species $j$. We can combine these species embeddings with a location embedding $f_{\bm{\theta}}(\bm{x})$ via an inner product to compute the probability that the species $j$ is present at $\bm{x}$. At inference time, we compute the presence probabilities for all species in the training set, for all locations $c$ in the set of available context locations $\mathcal{C}^{t}$ for our target species $t$. We then produce a new species embedding $\bm{w}_{t}$ by taking a weighted average of the existing $\bm{w}_{j}$’s where the weight for each is the product of the probabilities of presence for that species:
@@
We can then use this new species embedding for our target species to produce a probability of presence for any location $x$ as in SINR (Cole et al., 2023). We present few-shot results using this method in Figure 3. We see that the performance of the Active SINR approach is competitive with FS-SINR when provided with no text, though worse than FS-SINR when provided with this additional context. However, increasing the number of provided context locations beyond a small number actually hurts performance, as it is unable to accurately represent the range of a previously unseen species via the weighted combination of those from the training set.
C.3 Evaluation
@@ Appendix D Additional Ablations
Here we present additional results to investigate the impact of a variety of design choices and training procedures for FS-SINR. We present plots on a “Symlog” scale, where a linear scale is used between 0 and 10, in order to allow us to show zero-shot results alongside few-shot results. We display the mean of three runs with standard deviations shown as error bars and also show just the mean values alongside for easier interpretation.
D.1 Ablating Training Context Locations
@@
In Figure A12 we show ‘Range Text’ evaluation performance on the IUCN dataset for FS-SINR models trained using different amounts of context information at training time. We see that generally increasing the context used during training improves performance, and that having a fixed number of context locations is also beneficial.
@@
Figure A12: Impact of number of training context locations. Here we evaluate FS-SINR models trained using different numbers of context locations. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with ‘Range Text’ on the IUCN dataset. ‘Fixed’ indicates the same number of context locations were provided for every training example. ‘Variable’ indicates that a uniform random distribution of context locations up to the specified number were provided with each training example. We see that ‘Variable’ generally under-performs compared to ‘Fixed’ and that increasing the train context length tends to increase evaluation performance.
D.2 Ablating Context Information
@@
In Figure A13 we display ‘Range Text’ evaluation performance on the IUCN dataset for FS-SINR models trained using different combinations of text and context locations during training. We observe that good text-only zero-shot performance requires sometimes providing just text as context information during training. This forces the model to learn to produce ranges from only text information. Models that are sometimes provided with both text and locations for the same training examples perform best as the number of provided context locations increases. We also see that models trained without text can perform on par with those that see text during training when enough context locations are provided (5 - 10). As we might expect, models that are provided with token types they have not seen during training perform poorly.
@@
Figure A13: Impact of train context information. Here we evaluate FS-SINR models trained using different context information on the IUCN dataset. Results are shown with standard deviations from three runs (left), and without (right) for clarity. Evaluation is performed with ‘Range Text’ unless ‘No Eval Text’ is specified, in which case just locations are provided during eval. 70% of training examples for ‘Default FS-SINR’ provide both location and text context, 20% provide just locations 10% and provide just text. ‘Always Obs. Only’ has only seen locations during training. ‘Always Text Only’ has only seen Text during training. ‘Always Text and Obs’ is always provided with both locations and text during training. ‘Always Text or Obs.’ is provided with just locations for 90% of training examples, and just text for the remaining 10%.
D.3 Ablating Input Features
@@
In Table A5 we provide additional zero-shot results expanding on those in Table 1 from the main paper. Specifically, we add comparisons to using a different location encoder (i.e., SATCLIP (Klemmer et al., 2025) instead of SINR), comparisons to using a DINOv2 pre-trained image encoder (DINOv2-large) (Oquab et al., 2024), comparisons to using the environmental covariates as in SINR (Cole et al., 2023) that contain information about a locations’ climate in addition to the spatial coordinates.
@@
Table A5: Additional zero-shot results. We report zero-shot performance where no location information is provided to each model, comparing SINR (Cole et al., 2023), LE-SINR (Hamilton et al., 2024), and variants. We denote additional metadata as: EN for additional environmental covariates (Fick & Hijmans, 2017) used in Cole et al. (2023), HT for ‘Habitat Text’, RT for ‘Range Text’, I for ‘Image’ using our default EVA-02 image encoder, I (DINOV2) for Image using a DINOV2 based image encoder (Oquab et al., 2024), TST for ‘Test Species in Train’, TRT for using full taxonomic rank text, SATCLIP for where the SINR encoders are replaced with the image derived location encoders from Klemmer et al. (2025), and P-LE-SINR for ‘Prototype LE-SINR’. Results are reported as MAP, where higher is better.
@@
D.4 Ablating Location Encoder